OpenStudy (anonymous):

i need help with this please , looks simple but its hard 4(x + 1) = −4

OpenStudy (anonymous):

distribute out the 4 in the left hand side. What do you get then?

OpenStudy (anonymous):

x+1=-8?

OpenStudy (anonymous):

Do you know what distribute out means?

OpenStudy (anonymous):

i may have at one point but i dont remember

OpenStudy (anonymous):

4(x + 1) when I mean distribute I mean you have to carry multiplication of 4 to the inside of the parentheses. That means you multiply 4 to x and you also multiply 4 to 1

OpenStudy (anonymous):

ooh ok i remeber know , im slow , anyways , wont it be 4x+1 , how would that work?

OpenStudy (anonymous):

um close but you forgot to multiply 4 to 1 as well

OpenStudy (anonymous):

you should get 4x+4

OpenStudy (anonymous):

4(x + 1) = −4 would then be 4x + 4 = −4

OpenStudy (anonymous):

The next step is subtracting both sides by 4 because we want to get rid of that 4 in the left side of the equation

OpenStudy (anonymous):

ok so it would be 4=-8 wich is -2 right ?

OpenStudy (anonymous):

4x + 4 = −4 4x = −8 x = −8/4 x = −2

OpenStudy (anonymous):

\[x=-2\]

OpenStudy (anonymous):

yup you are right but you can't forget about the x. you have to write it down or the teacher might take some points off

OpenStudy (anonymous):

yeah so i had it right i just didnt put the x= part

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yup